# World Model on Million-Length Video And Language With Blockwise RingAttention

### v3

@@ 3.2 Training Steps
We initialize from LLaMA-2 7B [55] and progressively increase the effective context length of the model in 5 stages - 32K, 128K, 256K, 512K, and 1M. For each stage, we train on different filtered versions of the Books3 dataset from The Pile [19]. Table 1 details each information about each training stage, such as number of tokens, total time, and the Books3 dataset filtering constraints. Each successive run is initialized from the run of the prior sequence length.
3.3 Chat Fine-tuning for Long-Context Learning
Constructing QA data for Long Context Reasoning.  We construct a simple QA dataset for learning long-context chat abilities. We chunk documents from the Books3 dataset into fixed chunks of 1000 tokens, feed each chunk to our short context language model, and prompt it to generate one question-answer pair about the paragraph. Then, given a context length such as 32K, we construct a single 32K token example by concatenating adjacent chunks together, as well as appending relevant QA pairs towards the end of the sequence in chat form.
Training Details.  For chat fine-tuning, we train each model on a mix of UltraChat [14] and our custom QA dataset, with roughly a 7:3 ratio. We found it crucial to pre-pack the UltraChat data to the training sequence length, and keep them separate from examples with our QA data, as UltraChat data generally has a much higher proportion of loss tokens (densely packed, short chat sequences), whereas our QA data has a much lower percentage of loss tokens per sequence (< 1%) since there is no loss on the long documents that are in the given context. Table 2 shows further training details for each run. Note that progressive training is not performed very each of the chat models, and instead are initialized from their respective pretrained models at the same context length.
Figure 5: Needle retrieval task. Our LWM-Text-Chat-1M have near perfect accuracy across different positions in 1M context window.
@@ 3.3 Chat Fine-tuning for Long-Context Learning
3.4 Language Evaluation Results
3.4.1 Single Needle Retrieval
We evaluate on the popular Needle In A Haystack task [22] - more specifically an easier to evaluate version [2] that finds and retrieves random numbers assigned to randomized cities from the context. Figure 5 shows nearly perfect retrieval accuracy over the entire context of our 1M context model. In addition, Figure 2 shows that we can scale to far larger contexts compared to the current best available LLMs. Appendix A shows more single needle retrieval results for our other shorter context length models.
@@ 3.4.1 Single Needle Retrieval
3.4.2 Multi-Needle Retrieval
We additionally examine the performance of our model on more complex variant of the needle retrieval task by mixing in multiple needles, as well as trying to retrieve a specific subset of them. Figure 6 shows multi-needle retrieval results under different settings. Our model generalizes well when retrieving a single needle from multiple needles in context, with slight degradation when asked to retrieve more than one needle. Table 3 shows multi-needle comparisons between our model, Gemini Pro, and GPT-4, where our model is able to perform competitively or better than GPT-4 at retrieving one needle, or slightly lower performance when retrieving more than one needle. Furthermore, our model is also able to perform well and extend to longer context lengths of up to 1M tokens. However, we note that we see degradation in accuracy while increasing the difficulty of the needle retrieval task, suggesting that there is still more room to improve on the 1M context utilization of our model. We believe that our released model will provide a foundation for future work on developing longer context models, as well as encourage more challenging benchmarks that contain difficult long-range tasks that require higher levels of synthesis, rather than pure fact retrieval.
Table 3: Multi-Needle Retrieval Accuracy Baseline Comparison
Context Length Model $N=2,R=2$ $N=4,R=1$ $N=4,R=2$
32K Gemini Pro 0.34 0.44 0.6
GPT-4 0.97 0.95 0.9
LWM-Text-1M Ours) 0.84 0.97 0.84
128K Gemini Pro - - -
GPT-4 0.92 0.8 0.82
LWM-Text-1M (Ours) 0.83 0.98 0.83
1M Gemini Pro - - -
GPT-4 - - -
LWM-Text-1M (Ours) 0.67 0.84 0.69
3.4.3 Short Context Language Evaluation
Table 4: Evaluation of language tasks: Comparison between Llama2-7B (4K context) and context-expanded versions of LWM-Text: 32K to 1M. Results indicate that expanding context does not negatively impact performance on short-context tasks.
LWM-Text
Task / Metric Llama-2 7B 32k 128k 256k 512k 1M
arc_challenge/acc 0.4 0.43 0.45 0.44 0.44 0.43
arc_challenge/acc_norm 0.43 0.47 0.47 0.46 0.46 0.46
hellaswag/acc 0.57 0.57 0.57 0.57 0.56 0.57
hellaswag/acc_norm 0.77 0.76 0.76 0.76 0.75 0.75
mmlu 0.39 0.4 0.41 0.41 0.36 0.35
openbookqa/acc 0.32 0.33 0.31 0.32 0.33 0.3
openbookqa/acc_norm 0.44 0.44 0.44 0.43 0.41 0.41
3.4.4 Chat Evaluation
We additionally evaluate the our model on MT-Bench [60] to test its conversation ability. Table 6 shows the MT-Bench scores of for each of our models. Table 6 illustrates the relationship between the mix of chat and fact retrieval tasks and the performance on MT-Bench score and Needle Retrieval accuracy. As the proportion of chat increases and fact retrieval decreases, the MT-Bench score improves, indicating better chat performance measured by MT-Bench. Conversely, Needle Retrieval accuracy decreases, suggesting a trade-off where increasing chat interaction capabilities may reduce the system’s precision in retrieving specific information or ’needles’ from input context. Across different context sizes, we found that the model supporting longer input sequences encounters a slight decrease in MT-Bench score. We hypothesize that this is because we chose to train with fewer examples on longer sequence training and can be improved by simply training on more data. In addition, this trade-off may be resolved by acquiring higher quality long-context chat data that is closer to the chat distribution of the UltraChat dataset.
@@ 3.4.4 Chat Evaluation
4 Stage II: Learning Long-Context Vision-Language Models
4.1 Architectural Modifications For Vision
The model is illustrated in Figure 4. We use the pretrained VQGAN [17] from aMUSEd [45] that tokenizes $256\times 256$ input images to $16\times 16$ discrete tokens. Videos are tokenized by applying the VQGAN per-frame, and concatenating the codes together. In order to distinguish between modalities when generating, as well as knowing when to switch, we introduce mechanims to mark the end of text generation / beginning of vision generation, and vice-versa. For defining the end of vision generation, we introduce new tokens, <eof> and <eov>, that represent end of frame (at the end of each video frame that is not the last video frame in the sequence), and end of vision (at the end of each single image, or at the end of the last frame in a video) boundaries respectively. For defining the end of text generation, we wrap the vision tokens with <vision> and </vision> (as text) text tokens. The model is trained with interleaved concatenations of vision and text tokens, and predicted autoregressively.
Table 7: LWM and LWM-Chat Training Stages
1K 8K Chat-32K Chat-128K Chat-1M
Parameters 7B 7B 7B 7B 7B
Sequence Length $2^{10}$ $2^{13}$ $2^{15}$ $2^{17}$ $2^{20}$
RoPE $\theta$ 50M 50M 50M 50M 50M
Tokens per Batch 8M 8M 8M 8M 8M
Total Tokens 363B 107B 10B 3.5B 0.4B
Wall Clock 83h 32h 10h 6h 8h
Compute (TPU) v4-1024 v4-1024 v4-1024 v4-1024 v4-1024
4.2 Training Steps
  • LWM-1K: We train on large set of text-image dataset comprising of a mix of LAION-2B-en [51] and COYO-700M [9]. The datasets were filtered to only include images with at least 256 resolution – in total roughly 1B text-image pairs. During training, we concatenate the text-image pairs and randomly swap the order of the modalities to model both text-image generation, unconditional image generation, and image captioning. We pack text-image pairs to sequences of 1K tokens.
  • LWM-8K: We train on a text-video dataset mix of WebVid10M [7] and 3M InternVid10M [58] examples. Similar to prior works [25, 27, 57], we jointly train on both images and video with a 50-50 ratio of each modality. We pack images to sequences of 8K tokens, and 30 frame videos at 4FPS. Similar to image training, we randomly swap the order of modalities for each text-video pair.
  • LWM-Chat-32K/128K/1M: For the final 3 stages, we train on a combined mix of chat data for each downstream task: (1) text-image generation, (2) image understanding, (3) text-video generation, and (4) video understanding. We construct a simple version of text-image and text-video chat data by sampling random subsets of the pretraining data augmented with chat format. For image understanding, we using the image chat instruct data from ShareGPT4V [11]. Lastly, for the video understanding chat data, we use a combined mix of Valley-Instruct-73K [42] and Video-ChatGPT-100K instruct data [43]. For all short context data (image generation, image understanding, video generation), we pack sequences to the training context length. During packing, we found it crucial to mask out the attention so that each text-vision pair only attends to itself, as well as re-weighting losses to make computation identical to training in a non-packed + padding training regime. For video understanding data, we uniformly sample a max number of frames to fit the training context length of the model if the video is too long. During training, We allocate 25% of each batch to each of the 4 downstream tasks.
For the first two stages of training (LWM-1K and LWM-8K), we additionally mix 16% of the batch to be pure text data from OpenLLaMA [23], as we found it beneficial to preserve language capabilities while training on vision data.
((a)) The assistant answers questions about a 3-minute video of animals.
((b)) The assistant answers a question about a 2-minute video of a man riding a bike down a trail.
Figure 7: LWM can answer questions about videos. More examples can be found in Appendix B.
4.3 Vision-Language Evaluation Results
4.3.1 Long Video Understanding
Although vision-language model [41, 44, 52] can ingest long videos, this is commonly done by performing large temporal subsampling of video frames due to limited context length. For example, Video-LLaVA [41] is restricted to uniformly sampling 8 frames from a video, no matter how long the original video may be. As such, models may lose more fine-grained temporal information that is important for accurately answering any questions about the video. In contrast, our model is trained on long sequences of 1M tokens, and as a result, can simultaneously attend thousands of frames of videos to retrieve fine-grained information over short time intervals. Figure 1 shows an example of our model correctly answering questions about a long, 1-hour YouTube compilation consisting of more than 500 individual clips. Our baseline methods, on the other hand, generally have difficulty answering the questions due to a limited number of frames. More results are shown in Figure 7 and Appendix B.
Although we demonstrate that our model can perform QA over complex, long-form videos, we note that there is still room to improve for better context utilization across all 1M tokens, as generated answers from our model may not always accurate, and the model still struggles with more complex questions that require higher-level understanding of the video. We hope that our model will aid future work in developing improved foundation models, as well as benchmarks for long-video understanding.
4.3.2 Image Understanding and Short Video Understanding
Tables 8 and 9 show results on common benchmarks for image understanding and short video understanding. Figure 17 shows qualitative examples for image understanding. Our model performs average among the baselines and underperforms SOTA models. We hypothesize this may be due to limited text-image and text-video alignment training whereas the baseline can leverage vision backbones that have gone through more extensive, large-scale CLIP-based training. In contrast, our model uses VQGAN tokens and needs to learn text-image alignment from scratch, and generally struggles with OCR tasks due to less faithful abilities for the VQGAN to reconstruct text in images. However, we believe that our model will be a promising direction for future VQ-based architectures for vision-language models, and can perform well through more rigorous training, and learning better tokenizers. Appendix B shows more qualitative image understanding and Appendix B shows more qualitative video understanding examples.
Table 8: Image Understanding Benchmarks
{adjustwidth}
-0.2cm Method Visual Token VQAv2 GQA VisWiz SQA TextVQA POPE MM-Vet MiniGPT-4 [61] CLIP - 30.8 47.5 25.4 19.4 - 22.1 Otter [40] CLIP - 38.1 50 27.2 21.2 - 24.6 InstructBLIP [16] CLIP - 49.2 34.5 60.5 50.1 - 26.2 LLaVA-1.5 [33] CLIP 78.5 62 38.9 66.8 58.2 85.9 30.5 LWM (ours) VQGAN 55.8 44.8 11.6 47.7 18.8 75.2 9.6
Table 9: Video Understanding Benchmarks
{adjustwidth}
-0.2cm MSVD-QA MSRVTT-QA TGIF-QA Method Visual Token Accuracy Score Accuracy Score Accuracy Score VideoChat [32] CLIP 56.3 2.8 45 2.5 34.4 2.3 LLaMA-Adapter [21] CLIP 54.9 3.1 43.8 2.5 - - Video-LLaMA [63] CLIP 51.6 2.5 29.6 1.8 - - Video-ChatGPT [43] CLIP 64.9 3.3 49.3 2.8 51.4 3 Video-LLaVA [41] CLIP 70.7 3.9 59.2 3.5 70 4 LWM (ours) VQGAN 55.9 3.5 44.1 3.1 40.9 3.1
4.3.3 Image and Video Generation
In addition to being able to perform image / video captions, as well as QA, our model can also generate images and video from text. Figure 8 shows some examples of such capabilities. We use classifier-free guidance [26] on the logits for autoregressive sampling in a manner similar to prior work [59, 24]. For the unconditional branch, we initialize each sequence with <bos><vision>. Appendix E and D show more image and video generation examples.
Figure 8: LWM can generate images and videos given text input. Examples of image and video generations. More examples are shown in Appendix E and Appendix D.
4.3.4 Masked Sequence Packing Ablation.
Table 10: Ablation study comparing standard and masked sequence packing mechanisms across three tasks. Masked sequence packing is crucial for performance.
VQAv2 SQA POPE
Naive Packing 48.3 34.8 62.5
LWM (Ours) 55.8 47.7 75.2
5 Further Details
MFU.  We trained our models using TPUv4-1024, which is approximately equivalent to 450 A100s, with a batch size of 8M using FSDP [18], Blockwise Transformer, and RingAttention for large contexts. Figure 9 shows the model FLOPS utilization (MFU) for each training stage. Blue color bars show language training and orange color bars show vision-language training. Our training achieves good MFUs even for very large context sizes.
Training Loss Curves.  Figures 10 and 11 show the training loss curves for each stage of training the language and vision-language models respectively.
Figure 10: Train loss curve for each training stage for LWM-Text models.
Figure 11: Train loss curve for each training stage for LWM and LWM-Chat models. Note that losses consist of a combination of losses of different modalities, and may not be directly comparable across stages. The sharp peak in the middle of 1K training is due to newly incporating EOF and EOV tokens into the vision codebook.
Training Hyperparameters. See Appendix F
Scaling Inference We additionally scale our inference code to support million-length sequences by implementing RingAttention for decoding. Inference for such long sequences requires a minimum of v4-128 with a TPU mesh sharding of 32 tensor parallelism, and 4 sequence parallelism (ring dimension). We perform inference in pure single precision, where additional improvements can be made through techniques in scalability such as quantization.
6 Related Works
Our work is related to efforts that extend language models’ context windows to allow more tokens [13, 56, 38, inter alia], often by using novel extrapolation methods to expand pretrained positional encodings followed by finetuning the model on longer context data. Our model adopts a simple approach by gradually increasing $\theta$ in RoPE positional encodings along with the training context window sizes, and we found this method to be effective. There has been research into architectures that do not model pairwise interaction, such as sparse attention and sliding window [10, 8]. Parallelization along the sequence dimension is studied in [37], but is not optimized for blockwise transformers and is incompatible with memory-efficient attention, which are crucial to large context training. [30] propose hybrid usage of tensor parallelism and sequence parallelism to better reduce memory cost on a single device. Our work utilizes RingAttention and Blockwise Transformer [39, 31] to model exact pairwise interactions in very long sequences for optimal performance. Concurrent to our work, Gemini 1.5 [49] achieves millions of tokens context size. Further training performance improvement is also possible with the load balancing of skipping causal masked computation [6, 36].
Our work is also related to works on instruction tuning [53, 12, 20, inter alia]. These studies focus on finetuning models using conversational data to enhance their capability across language tasks. Our approach seeks to advance models’ understanding of complex, long sequences of videos and languages. To achieve this, we extend the models’ context size by training on books and long videos, and finetuning on model-generated QA data to learn chat ability over long sequences.
Our work is also related to efforts combining vision and language [34, 41, 1, 62, 29, 3, inter alia]. These efforts often use CLIP [47] or BLIP [35] to encode visual information into embeddings for inputting into language models. They have the potential advantages of leveraging CLIP’s cross-modal understanding for encoding textual information in images. However, they can only be trained to predict text given visual input but not vice versa, limiting their ability to learn from diverse formats of visual and language information. In cases where they do predict visual tokens [29], a stronger diffusion decoder is generally required due to the relatively lossy nature of CLIP embeddings. Our work, on the other hand, is autoregressive "tokens in, tokens out", allowing us to flexibly model diverse forms of image-text, text-image, text-video, video-text, and purely video, image, or text formats.
7 Conclusion
In this paper, we address the challenge of learning models to understand the world better by combining language and video. We utilize RingAttention and Blockwise Transformers to scalably train on a massive dataset of long videos and books and gradually increase sequence length, from 32K to 1M tokens, to keep compute manageable. We develop masked sequence packing and loss weighting to effectively train on a diverse dataset of videos, images, and books. Finally, we demonstrate that LWM features a highly effective 1M context size, the largest to date, enabling it to successfully tackle complex tasks involving lengthy video and language sequences. We open source our optimized implementation of Blockwise RingAttention, masked sequence packing and other key features for training on millions-length sequences, as well as a 7B parameter model capable of processing over 1M multimodal tokens. We hope this work paves the way for advancing AI models with a reliable reasoning and a grounded understanding of the world and broader capabilities.
Limitations and Future Work.
Although this work achieves an effective very large context of over 1M tokens for large autoregressive models, and shows promising results in understanding over 1-hour-long videos and long-form language sequences, it does have some limitations that need to be addressed:
  • Better Video Tokenization. This work uses image tokenizer for videos – improving tokenization to be more compact could not only enhance video quality but also enable the processing of significantly more complex and longer videos, allowing more capable world model.
  • More Modalities. Our work paves the road for learning from additional modality sources such as audio and other long sequences, allowing a deeper and more holistic understanding about the world.
  • Better and More Video Data. Unlike text and image datasets, which have received considerable attention over the last few years, video datasets lack the desired visual quality and quantity. Future research can address this by sourcing YouTube videos.
Acknowledgments
This project is supported in part by Office of Naval Research grant N00014-21-1-2769 and Multidisciplinary University Research Initiative (MURI) award by the Army Research Office (ARO) grant No. W911NF-23-1-0277. We thank Google TPU Research Cloud for granting us access to TPUs, and thank Google Cloud for granting us research credits for storage.
Appendix A More Single-Needle Retrieval Results
Figure 12: Single needle retrieval accuracy for LWM-Text-Chat-256K
Figure 13: Single needle retrieval accuracy for LWM-Text-Chat-512K
Appendix B More Video Understanding Examples
Figure 14: LWM demonstrates video understanding over 1 hour video.
Figure 15: LWM demonstrates video understanding over 1 hour video.

### v4

@@ 4.2 Training Steps
  • LWM-1K: We train on large set of text-image dataset comprising of a mix of LAION-2B-en (Schuhmann et al., 2022) and COYO-700M (Byeon et al., 2022). The datasets were filtered to only include images with at least 256 resolution – in total roughly 1B text-image pairs. During training, we concatenate the text-image pairs and randomly swap the order of the modalities to model both text-image generation, unconditional image generation, and image captioning. We pack text-image pairs to sequences of 1K tokens.
  • LWM-8K: We train on a text-video dataset mix of WebVid10M (Bain et al., 2021) and 3M InternVid10M (Wang et al., 2023) examples. Similar to prior works (Ho et al., 2022a; b, Villegas et al., 2022), we jointly train on both images and video with a 50-50 ratio of each modality. We pack images to sequences of 8K tokens, and 30 frame videos at 4FPS. Similar to image training, we randomly swap the order of modalities for each text-video pair.
  • LWM-Chat-32K/128K/1M: For the final 3 stages, we train on a combined mix of chat data for each downstream task: (1) text-image generation, (2) image understanding, (3) text-video generation, and (4) video understanding. We construct a simple version of text-image and text-video chat data by sampling random subsets of the pretraining data augmented with chat format. For image understanding, we using the image chat instruct data from ShareGPT4V (Chen et al., 2023a). Lastly, for the video understanding chat data, we use a combined mix of Valley-Instruct-73K (Luo et al., 2023) and Video-ChatGPT-100K instruct data (Maaz et al., 2023). For all short context data (image generation, image understanding, video generation), we pack sequences to the training context length. During packing, we found it crucial to mask out the attention so that each text-vision pair only attends to itself, as well as re-weighting losses to make computation identical to training in a non-packed + padding training regime. For video understanding data, we uniformly sample a max number of frames to fit the training context length of the model if the video is too long. During training, We allocate 25% of each batch to each of the 4 downstream tasks.
Figure 6: LWM excels in answering questions about a 1-hour YouTube video. This figure compares LWM-Chat-1M with proprietary models like Gemini Pro Vision and GPT-4V, along with open-source models. The test involves answering questions based on an hour-long YouTube compilation containing over 500 video clips. LWM demonstrates superior performance in providing accurate answers requiring comprehension of extended video content.
For the first two stages of training (LWM-1K and LWM-8K), we additionally mix 16% of the batch to be pure text data from OpenLLaMA (Geng and Liu, 2023), as we found it beneficial to preserve language capabilities while training on vision data.
Summary: Stage II training incorporates image and video. Building on Stage I, it gradually increases sequence lengths of vision and text input. Importantly, we found our masked sequence packing and mixing synthetic and chat data crucial to retain short context performance during our progressive training. Appendix B shows ablations when not using our training method on instruction-following and text-image understanding benchmarks.
4.3 Vision-Language Evaluation Results
4.3.1 Long Video Understanding
Although vision-language model (Lin et al., 2023, OpenAI, 2023, Team et al., 2023) can ingest long videos, this is commonly done by performing large temporal subsampling of video frames due to limited context length. For example, Video-LLaVA (Lin et al., 2023) is restricted to uniformly sampling 8 frames from a video, no matter how long the original video may be. As such, models may lose more fine-grained temporal information that is important for accurately answering any questions about the video. In contrast, our model is trained on long sequences of 1M tokens, and as a result, can simultaneously attend thousands of frames of videos to retrieve fine-grained information over short time intervals. Table 4 shows long video evaluations on the Video-MME (Fu et al., 2024) benchmark, demonstrating our model as the best performing model among its size class. Figure 6 shows an example of our model correctly answering questions about a long, 1-hour YouTube compilation consisting of more than 500 individual clips. Our baseline methods, on the other hand, generally have difficulty answering the questions due to a limited number of frames. More results are shown in Figure 18 and Appendix F.
Table 4: Long Video-MME Benchmark. * denotes models after the completion of this paper.
Method Parameters Frames Medium (4min-15min) Long (30min-60min)
Gemini 1.5 Pro* Unknown $\leq 1800$ $74.3$ $67.4$
GPT-4o* Unknown $384$ $70.3$ $65.3$
LLaVA-Video* 72B $64$ $68.9$ $61.5$
VideoLLaMA 2* 72B $32$ $59.9$ $57.6$
Long-LLaVA* 7B $64$ $51.4$ $45.4$
Video-LLaVA 7B $8$ $38.1$ $36.2$
LWM-1M 7B $\leq 1800$ $63.7$ $60.8$
Table 5: Image Understanding Benchmarks (left) and Video Understanding Benchmarks (right)
Method Visual Token VQAv2 GQA SQA MiniGPT-4 CLIP - 30.8 25.4 Otter CLIP - 38.1 27.2 InstructBLIP CLIP - 49.2 60.5 LLaVA-1.5 CLIP 78.5 62.0 66.8 LWM (ours) VQGAN 55.8 44.8 47.7       Method MSVD MSRVTT TGIF VideoChat 56.3 45 34.4 LLaMA-Adapte 54.9 43.8 - Video-LLaMA 51.6 29.6 - Video-ChatGPT 64.9 49.3 51.4 LWM (ours) 55.9 44.1 40.9
4.3.2 Image Understanding and Short Video Understanding
We evaluate LWM on standard benchmarks for image and short video understanding, with results presented in Table 5. Our model performs comparably to baselines but falls short of state-of-the-art (SOTA) models. This performance gap is not unexpected, given that SOTA models leverage vision backbones that have undergone extensive CLIP training (Radford et al., 2021). In contrast, LWM utilizes discrete tokens from an off-the-shelf model (Patil et al., 2024). Discrete tokens result in greater information loss, particularly for OCR-like textual data, compared to continuous CLIP embeddings. Moreover, our model learns text-image alignment from scratch, while CLIP-based models benefit from large-scale pretraining. This work primarily focuses on long-context methodology, and we defer additional training to future work due to computational constraints. A straightforward approach to improving benchmark scores would be to incorporate CLIP embeddings as additional input. Despite not achieving SOTA scores on these short video benchmarks, we believe LWM provides valuable insights for future long-context language and video understanding and generation. The model’s performance could be enhanced through additional training and minor modifications. We include qualitative image understanding examples in Appendix E and qualitative video understanding examples in Appendix F.
4.3.3 Image and Video Generation
Thanks to a unified any-to-any architecture, our model can not only perform image/video captioning and question-answering but also generate images and videos from text. Figure 7 demonstrates examples of these capabilities. For autoregressive sampling, we employ classifier-free guidance (Ho and Salimans, 2022) on the logits, similar to previous works (Yu et al., 2022, Gafni et al., 2022). In the unconditional branch, we initialize each sequence with <bos><vision>. For additional image and video generation examples, please refer to Appendices H and I, respectively.
Takeaway: LWM excels in long video understanding by processing significantly more frames than previous state-of-the-arts, resulting in better understanding. Moreover, its long-context enabled unified any-to-any architecture allows for versatile image and video and text understanding and generation.
Figure 7: LWM’s ability to generate both static images and dynamic videos from text is shown. The top row illustrates image, while the bottom rows show video.
5 Related Works
Our research builds upon existing efforts to extend the context windows of language models, enabling them to process more tokens (Chen et al., 2023b, Tworkowski et al., 2023, Liu et al., 2023c). These approaches often employ innovative extrapolation techniques to expand pretrained positional encodings, followed by model finetuning on longer context data. In contrast, our model takes a straightforward approach by incrementally increasing $\theta$ in RoPE positional encodings alongside expanding the training context window sizes, which we found to be effective. Additionally, there have been investigations into architectures that avoid modeling pairwise interactions, such as sparse attention and sliding window techniques (Child et al., 2019, Beltagy et al., 2020). Prior research has explored sequence parallelization (Li et al., 2021, Korthikanti et al., 2022, inter alia), though it is not optimized for blockwise transformers or compatible with memory-efficient attention, both of which are critical for large context training. Our work further leverages large context transformer techniques (Liu et al., 2024, Liu and Abbeel, 2023) to capture exact pairwise interactions in extended sequences for enhanced performance. Load-balancing strategies, such as skipping causal masked computation (Brandon et al., 2023, Li et al., 2023) offer room for further optimization. Concurrent developments like Gemini 1.5 (Reid et al., 2024) reach 1M tokens context size in language and video.
Additionally, our approach relates closely to advances in instruction tuning (Taori et al., 2023, Chiang et al., 2023, Geng et al., 2023), which focus on finetuning models with conversational data to boost their performance across diverse language tasks. We aim to extend these capabilities to the domain of long-sequence understanding in both video and language tasks. To achieve this, we extend the model’s context size by training on comprehensive datasets, including books and long videos, and finetune on model-generated question-answering datasets to enhance its ability to handle extended conversational sequences.
Furthermore, our research draws from work on integrating vision capabilities into language models (Liu et al., 2023b, Lin et al., 2023, Awadalla et al., 2023, Zhang et al., 2023, Jin et al., 2023b, Aiello et al., 2023). These efforts frequently utilize continuous embeddings (Radford et al., 2021, Li et al., 2022) to encode visual information into embeddings for inputting into language models. While these approaches benefit from CLIP’s cross-modal understanding to encode textual information from images, their ability to predict text from visual input is limited, as is their capacity to learn from diverse visual-language formats. In contrast, our autoregressive model, which processes "tokens in, tokens out," allows greater flexibility in modeling various formats, including image-text, text-image, text-video, video-text, and pure formats like video, image, or text. Our method is compatible with these prior works, making it an interesting future direction to combine continuous embeddings as input with discrete tokens and a long-context autoregressive model.
6 Conclusion
In conclusion, this paper tackles the critical challenge of enabling long-context understanding in sequence models, which is vital for developing generally intelligent systems capable of processing large temporal sequences. By exploring the development of 1M context language and video-language models, the work sets new benchmarks in language retrieval and long video understanding. We outline approaches to data curation and progressive context extension, accompanied by an efficient open-source implementation for scalable training on long sequences. Moreover, we open-source a family of 7B parameter models capable of handling over 1M tokens in text and video.
Limitations. While this work successfully develop a large large context of over 1M text and video tokens, and demonstrate promising results in processing hour-long videos and long documents, there are still some limitations that need to be addressed:
  • Improved tokenization and embedding. This work uses a vanilla image tokenizer for images and frame-by-frame tokenization for videos. Future work could explore video tokenization that takes time redundancy into account, as well as including continuous embeddings as input to enrich image understanding.
  • Limited scale. Our models use more tokens per parameter than Chinchilla’s recommendation, but being much smaller than current large language models (100B+ parameters), our findings may not directly apply to them. Extrapolating to larger scales should be done cautiously, as different scaling behaviors could emerge at those larger sizes.
Acknowledgments
This project is supported in part by Office of Naval Research grant N00014-21-1-2769 and ARO MURI (2023) on Neuro-Inspired Distributed Deep Learning. We thank Google TPU Research Cloud for granting us access to TPUs, and thank Google Cloud for granting us research credits for storage. Pieter Abbeel holds concurrent appointments as a Professor at UC Berkeley and as an Amazon Scholar. This paper describes work performed at UC Berkeley and is not associated with Amazon.
Appendix A Further Details
Model Flops Utilization. We trained our models using TPUv4-1024, which is approximately equivalent to 450 A100s, with a batch size of 8M using FSDP (Facebook, 2023) and BlockwiseRingAttention (Liu et al., 2024) for large contexts. Figure 8 shows the model FLOPS utilization (MFU) for each training stage. Blue color bars show language training and orange color bars show vision-language training. Our training achieves good MFUs even for very large context sizes.
Training Loss Curves. Figure 9 and Figure 10 show the training loss curves for each stage of training the language and vision-language models respectively.
Figure 9: Training progress over multiple days for LWM-Text. Train loss curve for each training stage for LWM-Text models.
Figure 10: Training progress over multiple days for LWM. Train loss curve for each training stage for LWM and LWM-Chat models. Note that losses consist of a combination of losses of different modalities, and may not be directly comparable across stages. The sharp peak in the middle of 1K training is due to newly incporating EOF and EOV tokens into the vision codebook.
Training Hyperparameters. See Appendix J
Scaling Inference. We additionally scale our inference code to support million-length sequences by implementing RingAttention for decoding. Inference for such long sequences requires a minimum of v4-128 with a TPU mesh sharding of 32 tensor parallelism, and 4 sequence parallelism (ring dimension). We perform inference in pure single precision, where additional improvements can be made through techniques in scalability such as quantization.
@@ Appendix A Further Details
Table 8: LWM and LWM-Chat Training Stages
1K 8K Chat-32K Chat-128K Chat-1M
Parameters 7B 7B 7B 7B 7B
Sequence Length $2^{10}$ $2^{13}$ $2^{15}$ $2^{17}$ $2^{20}$
RoPE $\theta$ 50M 50M 50M 50M 50M
Tokens per Batch 8M 8M 8M 8M 8M
Total Tokens 363B 107B 10B 3.5B 0.4B
Wall Clock 83h 32h 10h 6h 8h
Compute (TPU) v4-1024 v4-1024 v4-1024 v4-1024 v4-1024
Appendix B Ablation Studies
B.1 Masked Sequence Packing
Table 9: Ablation study comparing standard independent packing and our masked sequence packing mechanisms across three tasks. Results show that masked sequence packing significantly improves performance across all tasks.
VQAv2 SQA POPE
Standard independent packing 48.3 34.8 62.5
Masked sequence packing (Ours) 55.8 47.7 75.2
B.2 Mixing Synthetic And Chat Data
We additionally evaluate the our model on MT-Bench (Zheng et al., 2023) to test its conversation ability. Table 11 shows the MT-Bench scores of for each of our models. Table 11 illustrates the relationship between the mix of chat and fact retrieval tasks and the performance on MT-Bench score and Needle Retrieval accuracy. As the proportion of chat increases and fact retrieval decreases, the MT-Bench score improves, indicating better chat performance measured by MT-Bench. Conversely, Needle Retrieval accuracy decreases, suggesting a trade-off where increasing chat interaction capabilities may reduce the system’s precision in retrieving specific information or ’needles’ from input context. Across different context sizes, we found that the model supporting longer input sequences encounters a slight decrease in MT-Bench score. We hypothesize that this is because we chose to train with fewer examples on longer sequence training and can be improved by simply training on more data. In addition, this trade-off may be resolved by acquiring higher quality long-context chat data that is closer to the chat distribution of the UltraChat dataset.
@@ B.2 Mixing Synthetic And Chat Data
Appendix C More Single-Needle Retrieval Results
Figure 11: Needle retrieval task using the LWM-Text-Chat-1M model. The model demonstrates near-perfect retrieval accuracy across various positions within the 1M context window, as reflected by consistently high scores at different depth percentages and context lengths.
Figure 12: Single needle retrieval accuracy for the LWM-Text-Chat-256K model. The model achieves near-perfect retrieval performance across various positions in the 256K context window, as shown by consistently high scores across all depth percentages and context lengths.
Figure 13: Single needle retrieval accuracy for the LWM-Text-Chat-512K model. The model demonstrates near-perfect retrieval performance across different positions within the 512K context window, as indicated by consistently high scores at varying depth percentages and context lengths.
Appendix D Details Of The Multi-needle Evaluations
We built on top of the original Needle in a Haystack problem (see Section 3.3.2 for details), which was to retrieve a single magic number randomly inserted into some long context. In this case, magic numbers are associated with random cities (“The magic number for San Francisco is 2521233” $\rightarrow$ “What is the magic number for San Francisco?”). We extend the task by introducing two variables N and R, where N is the number of needles (magic numbers + different cities) randomly inserted into the context, and R is the random subset of magic numbers asked to retrieve (“What are the magic numbers for San Francisco, …, and Bangkok?”). Correctness is computed by extracting out the numbers retrieved for each cities and checked with string matching.
Appendix E More Image Understanding Examples
Figure 14: Question answering based on image input using LWM. The assistant provides detailed responses to questions about capturing the essence of a serene lakeside scene and interpreting a creative twist on a classic artwork featuring a dog.
Appendix F More Video Understanding Examples
Figure 15: LWM demonstrates video understanding over a 1-hour video. The figure shows a comparison of different AI systems responding to a question about the number of lemons in a person’s car. While GPT-4V, Gemini Pro Vision, and Video-LLAVA fail to provide a correct response, LWM accurately identifies that there are three lemons in the car.
Figure 16: LWM demonstrates video understanding over a 1-hour video. The figure compares AI systems’ responses to the question ’What animal was standing on a piano?’ While GPT-4V, Gemini Pro Vision, and Video-LLAVA provide incorrect or incomplete answers, LWM correctly identifies that the animal standing on the piano was a cat.